# PALM: Pushing Adaptive Learning Rate Mechanisms for Continual Test-Time Adaptation

### v2

@@ Abstract
Real-world vision models in dynamic environments face rapid shifts in domain distributions, leading to decreased recognition performance. Using unlabeled test data, continual test-time adaptation (CTTA) directly adjusts a pre-trained source discriminative model to these changing domains. A highly effective CTTA method involves applying layer-wise adaptive learning rates for selectively adapting pre-trained layers. However, it suffers from the poor estimation of domain shift and the inaccuracies arising from the pseudo-labels. This work aims to overcome these limitations by identifying layers for adaptation via quantifying model prediction uncertainty without relying on pseudo-labels. We utilize the magnitude of gradients as a metric, calculated by backpropagating the KL divergence between the softmax output and a uniform distribution, to select layers for further adaptation. Subsequently, for the parameters exclusively belonging to these selected layers, with the remaining ones frozen, we evaluate their sensitivity to approximate the domain shift and adjust their learning rates accordingly. We conduct extensive image classification experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C, demonstrating the superior efficacy of our method compared to prior approaches.
@@ 1 Introduction
In a real-world setting, machine-perception systems [1] function in a rapidly changing environment. In such contexts, pre-trained vision models are susceptible to performance degradation caused by potential distribution shifts. For instance, while these models may have been trained on clean images, images captured by sensors could be corrupted by various weather conditions.
In the realm of addressing domain shifts due to image corruptions, test-time adaptation (TTA) has been gaining a lot of traction lately [5, 36]. Here, a pre-trained source model, with no access to its training (source) data due to privacy, is adapted to unlabeled test data through online training. One strategy in TTA involves mitigating domain shifts by adjusting the source model parameters using pseudo labels and minimizing the entropy of the model predictions [36]. While this approach effectively adapts the model to a single type of corruption, it faces two constraints when continually applied across a sequence of test tasks featuring different corruptions. Firstly, there is the issue of error accumulation since pseudo labels are generated for unlabeled test data, errors accumulate over time. Secondly, catastrophic forgetting occurs, resulting in a long-term loss of the model’s source knowledge due to adaptations and large parameter updates across tasks.
Continual test-time adaptation (CTTA) has been recently introduced to tackle these challenges. In [23, 6, 36, 28], stable optimization is achieved by updating the batch normalization (BN) parameters or re-formulating its statistics. CoTTA [37] includes full model updates, with a teacher-student network setup, to handle the above issues. To further improve the performance of CoTTA, EcoTTA [33] introduces the attachment of “meta networks", i.e., a BN layer and a convolution block, to the frozen source model for efficient adaptation.
More recently, the authors in [15] demonstrated that based on the type of distributional shift of the test data, a certain subset of layers of a pre-trained model can be “surgically" or heuristically chosen to fine-tune. Inspired by this finding, Layer-wise Auto-Weighting (LAW) [24] was introduced recently. LAW utilizes the Fisher Information Matrix (FIM) [27] to estimate domain shifts, automatically weighting trainable parameters and adjusting their learning rates (LRs) accordingly. While LAW achieves state-of-the-art results for CTTA, it still encounters two significant limitations: 1) It presupposes that layer-wise importance can be accurately estimated via FIM with initially generated pseudo-labels. Nevertheless, it is widely recognized that pseudo-labels are inherently noisy and unreliable. 2) Relying on directly accumulating such approximated importance from noisy labels to calculate domain shifts can lead to inaccuracies.
@@ 2 Related Works
Test-time adaptation (TTA): The goal of TTA is to adapt source models to unlabelled test data in real-time using a source-free online approach [12, 34, 36]. TENT [36] optimizes BN parameters by minimizing the Shannon entropy [31] of pseudo labels. BN Stats Adapt (BN-1) [28] replaces the BN statistics with those from corrupted test data, while AdaContrast [3] uses contrastive learning and a memory module to update the entire model. However, these methods address only single domain shifts and overlook continual shifts.
Continual TTA (CTTA): CTTA requires models to adapt to dynamic domain shifts. CoTTA [37] uses a teacher-student setup with full model updates. The teacher model is updated using an exponential moving average (EMA) to deal with catastrophic forgetting and error accumulation with time. EcoTTA [33] adds “meta networks" to a frozen source model, computing an $L_{1}$ regularization loss to prevent catastrophic forgetting. DSS [38] introduces a dynamic thresholding framework to differentiate between low and high sample quality.
Efficient model fine-tuning: Freezing model parameters to learn information from downstream data has witnessed a good body of work [30, 17, 45, 7]. Depending upon the distribution shift of the target task, in [15], the authors demonstrate that heuristically selecting a subset of convolutional blocks for training while freezing the remaining blocks works better than full fine-tuning.
Adaptive LR methods: Research such as [40, 43] demonstrates that deeper layers of a network extract task-specific features, unlike shallow layers. AutoLR [26] developed an adaptive LR scheme based on weight variations between layers, although these rates tend to mostly stay constant during training. For CTTA, LAW [24] uses the FIM [27] to assess domain shifts and adjust parameter LRs accordingly. However, using pseudo-labels to calculate FIM may lead to inaccuracies, and accumulating FIMs across domains can result in errors. Moreover, the adjustment of LR weights for each layer is constrained by the base LR, disregarding the relative “sensitivity" of parameters to the adaptation needs of the current batch and domain.
Parameter sensitivity: Neural network pruning literature extensively explores parameter sensitivity [14, 18, 35, 20]. It suggests that less sensitive parameters require higher LRs for effective adaptation, and vice-versa. Additionally, [16] introduces a concept of temporal sensitivity variation, influenced by uncertainty, as a factor in training models for tasks like natural language understanding and machine translation.
@@ 3 Preliminary
3.1 Problem Setting
CTTA aims to adapt a discriminative source model, parameterized as $\theta^{s}$ and trained on the source dataset $(\mathcal{X}^{\text{source}},\mathcal{Y}^{\text{source}})$, to multiple incoming tasks $\mathcal{T}_{i}=\{x_{k}\}_{k=1}^{K}$, at test-time. Each $\mathcal{T}_{i}$ signifies the $i^{th}$ task with $K$ batches, across varying data distributions. At time step t=0, the model is initialized to ${\theta^{s}}$. It is then gradually adapted to each incoming batch $x_{k}$ of $\mathcal{T}_{i}$ in an online manner, where the model parameters are updated to ${\theta^{t}}$. The source dataset is unavailable in this adaptation process due to privacy and storage constraints. With the continuum of tasks, at each $t$, the complete model parameters $\theta^{t}$ consist of a deep tensor with $N$ layers, expressed as $\theta^{t}=\{\theta^{t}_{1},\ldots,\theta^{t}_{N}\}$. Each layer $\theta^{t}_{n}$ can represented by $\theta^{t}_{n}=\{\theta^{t}_{1,n},\ldots,\theta^{t}_{j,n},\ldots\}$, where $\theta_{j,n}^{t}$ is an arbitrary parameter of layer $\theta_{n}^{t}$.
3.2 Core Motivations
We provide a detailed view of surgical fine-tuning [15], AutoLR [26], and LAW [24]. It is critical to understand the design principles and reasons for having an adaptive LR scheme for CTTA, which motivates this work.
@@ 3.2 Core Motivations
Inspired by [40, 43], AutoLR, proposed for supervised model fine-tuning on a downstream task, claims that the weight changes across layers should be in an increasing order, where the weight change $\delta w_{n}$ in the $n^{th}$ layer with weights $w_{n}$ at training epoch e is $\delta w_{n}=\|w_{n}^{e}\ -w_{n}^{(e-1)}\|$. They hypothesize that initial layers, which extract general features, should have minimal $\delta w_{n}$, while deeper layers extract task-specific features. An adaptive LR optimization scheme constrains weight variations to increase across layers, without assessing each layer’s importance during adaptation. However, the LRs for each layer show minimal deviation and remain consistent throughout the training process.
As the first adaptive LR work for CTTA, LAW hypothesizes that identifying and controlling the adaptation rate of the different layers is useful [15]. To understand the loss landscape and layer importance, the authors compute the gradient from the log-likelihood as $s(\theta_{j,n}^{t};x_{k})=\nabla_{\theta_{j,n}^{t}}\log(p(x_{k}))$, where $p(x_{k})$ is the predicted output. Then, the FIM [27], as a domain-shift indicator, is computed as a first-order approximation of the Hessian matrix. This layer-wise importance scheme completely relies on the pseudo-labels to compute the log-likelihood and is highly unreliable and uncertain. Due to the distribution shift, the label quality would begin to decrease for a continuum of tasks. The authors also propose to capture “domain-level" FIM ($\hat{F_{n}^{t}}$), for each layer, by a direct accumulation of the current layer-wise FIM (${F_{n}^{t}}$) as, $\hat{F}_{n}^{t}=\hat{F}_{n}^{t-1}+F_{n}^{t}$.
This means that LAW puts equal emphasis on $\hat{F}_{n}^{t}$ and ${F_{n}^{t}}$, for each layer. An issue arises when ${F}_{n}^{t}$ of the current batch is small but gets outweighed by a “relatively" larger $\hat{F}_{n}^{t}$. This raises questions on the amount of adaptation required by the $n^{th}$ layer for batch $x_{k}$ as the FIMs dynamically change for each test batch. This leads to a possible imbalance between the accumulated FIMs $\hat{F}_{n}^{t}$ and the current FIM ${F_{n}^{t}}$.
@@
4 Proposed Method
Figure 1: PALM framework: At time t, input batch $x_{k}$ is processed by the model, parameterized by $\theta^{t}$. The KL-divergence between the softmax predictions and a uniform distribution is backpropagated to select the layers with the gradient norm $\leq$ $\eta$, to quantify the uncertainty. The parameter sensitivities of these layers, as an indicator of domain shift meters, are computed to update their learning rates (Eq. 7). Finally, with the optimization objective in Eq. 11, we update the model with the adjusted learning rates of the parameters.
In this section, we introduce and formalize our proposed method for CTTA to address the key limitations of the works discussed above. Fig. 1 diagrammatically illustrates our entire approach.
4.1 Layer Selection by Quantifying Prediction Uncertainty
CTTA maximizes model performance in the current domain and so it is vital to assess prediction certainty amidst rapid domain shifts. In the literature on continual learning and TTA, several works utilize FIM [13, 29, 24]. However, the unavailability of true labels, with an unknown distribution, pushes us to capture the model prediction uncertainty and adapt selected layers, automatically chosen, to a batch of test data, unlike surgical fine-tuning [15].
Motivated by [11], which proposed an uncertainty scoring function from the gradient space, we compute the gradients of each layer. In essence, we backpropagate the KL divergence between the softmax probabilities and a uniform distribution u = [$\frac{1}{C}$, …, $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where $C$ is the total number of classes of the current task. This is a measurement of how far the distribution of the model’s prediction is from a uniform distribution. Intuitively, the KL divergence will be small with a large distribution shift. Let $h(\cdot)$ denote the classifier of the model. The loss $\mathcal{L}(\theta^{t})$, to compute the gradients, is computed as follows,
@@ 4.1 Layer Selection by Quantifying Prediction Uncertainty
4.2 Parameter Sensitivity as a Domain Shift Indicator
New studies have found that model parameters demonstrate levels of criticality [2, 44] across different positions. Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors [22, 25]. Inspired by this observation, we assert that parameters contribute differently during CTTA. From the previous stage, we automatically select the layers based on the prediction uncertainty. For parameters of these selected layers only, we compute their sensitivity i.e., an approximation of an error when it is removed. Motivated by [14, 21, 20], we compute the sensitivity of a parameter $\theta_{j,n}^{t}$ based on the loss (Eq. 1) as follows,
@@ 4.2 Parameter Sensitivity as a Domain Shift Indicator
Precisely, $S_{j,n}^{t}$ approximates the change in the loss landscape at $\theta^{t}$ when $\theta_{j,n}^{t}$ is removed. Sensitivity, computed as a “weighting factor" of model weights [21, 20, 16], determines the level of adaptation needed for model parameters. We utilize the gradients of parameters from the selected layers, computed by Eq. 2, to compactly derive Eq. 3 from the first-order Taylor series expansion of $\mathcal{L}(\theta^{t})$ [lee2018snip] to,
@@
Our rationale is as follows: parameters with low sensitivity are pushed to have higher LRs owing to more adaptation. However, $S_{j,n}^{t}$ of $\theta_{j,n}^{t}$ can be highly uncertain and would have a large variation. We introduce a method to capture this “domain-level" uncertainty by computing a weighted moving average of the sensitivity $S_{j,n}^{t}$. This is formulated as,
@@
where $\hat{S}_{j,n}^{t}$ denotes the “domain-level" sensitivity, computed on-the-fly, and $\alpha$in [0,1] is a smoothing factor. Specifically, $\alpha$ decides the contribution of the current sensitivity $S_{j,n}^{t}$ to $\hat{S}_{j,n}^{t}$. Since CTTA happens at the batch level, it is critical to achieve a good balance between $S_{j,n}^{t}$ and $\hat{S}_{j,n}^{t}$, with a slightly larger focus on $S_{j,n}^{t}$. As can be seen, we do not directly accumulate the “domain-level" information, unlike LAW. Due to large distribution shifts happening because of the continuum of tasks, larger deviations of $S_{n}^{t}$ from $\hat{S}_{n}^{t}$ are bound to happen. We quantify this statistic as an absolute difference,
@@
We emphasize that since Eq. 5 leverages sensitivity from previously computed batches and domains, $\hat{D}_{j,n}^{t}$ would act as a good indicator of how uncertain $S_{j,n}^{t}$ is to the previous batches and domains i.e. $\hat{S}_{j,n}^{t}$, for $\theta_{j,n}^{t}$. This would take care of error accumulation too. To automatically alter the LRs of these parameters, the “importance" ($i_{j,n}^{t}$) is calculated as,
@@
4.3 Optimization Objective
Though our method is robust, it is necessary to avoid miscalibrated labels. TENT [36] minimized the entropy on the output logits, l, following $\mathcal{H}(l)=-\sum_{c}p(l_{c})\log(p(l_{c}))$, where $p(l_{c})$ is the probability for class c. Following EcoTTA [33], we use entropy minimization as the objective for adaptation where $\mathcal{H}(l)$ is minimized for samples below a certain fixed threshold $\mathcal{H}_{0}$ as,
@@ 4.3 Optimization Objective
$\mathcal{H}_{0}$ is set to $0.4\times\log C$ following EcoTTA. For a fair comparison with prior works [37, 24], we apply a consistency loss between l and $\hat{l}$, computed as the output logits from the augmented version of $x_{k}$, to regularize the model as follows,
@@
5 Experiments
We draw a contrast of our proposed method against other state-of-the-art works in a continual test-time adaptation (CTTA) setting. However, to have a broader study and analysis, we also report results in a gradual test-time adaptation (GTTA) setting, following [37, 24]. In the following subsections, we describe the implementation details of both the challenging adaptation settings.
5.1 Experimental Settings and Baselines
Datasets and tasks. Following the standard benchmarks set by CoTTA [37], we evaluate our proposed method on CIFAR-10C, CIFAR-100C, and ImageNet-C, based on image corruption schemes as set in [9]. Each dataset contains a set of 15 corruption styles as tasks (e.g. gaussian noise, shot noise, impulse noise, $\ldots$) with 5 severity levels. We follow CoTTA and apply the same test-time augmentations, like color jitter, gaussian blur, gaussian noise, random flips, and random affine, to the input $x_{k}$. As the evaluation metric, we report the mean classification error.
CTTA implementation details. We evaluate our approach following prior state-of-the-art methods [36, 37, 3, 33, 28, 24]. To maintain fairness regarding the usage of pre-trained source models, we employ WideResNet-28 [42] for CIFAR-10C, ResNeXt-29 [39] for CIFAR-100C, and ResNet-50 [8] for ImageNet-C, all available on RobustBench [4]. We optimize using an Adam optimizer, setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C. For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively. We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout. Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA, and results are only reported for the highest severity level of 5 for each task.
GTTA implementation details. In the standard CTTA setup, each task only receives images with the highest severity. However, in GTTA, following CoTTA, for each task, the severity levels are gradually changed as $\ldots$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5$\rightarrow$4$\rightarrow$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\ldots$. We adhere to the details specified in the CTTA setting.
Baselines. We compare against several CTTA baselines, including BN Stats Adapt (BN-1) [28], TENT Continual [36], CoTTA [37], EcoTTA [33], AdaContrast [3], DSS [38], and LAW [24]. We also contrast our work against surgical fine-tuning [15]. For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 [39] for CIFAR-100C. Following the findings of surgical fine-tuning [15], we continually fine-tune only the first convolutional block of the respective models, denoted as “Surgical fine-tuning (continual)", using an SGD optimizer with fixed learning rates of 1e-3 for the CIFAR datasets and 5e-4 for ImageNet-C. For the reported results of LAW, we reproduce their results strictly following their official implementation [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.

### v3

@@ Abstract
Real-world vision models in dynamic environments face rapid shifts in domain distributions, leading to decreased recognition performance. Using unlabeled test data, continuous test-time adaptation (CTTA) directly adjusts a pre-trained source discriminative model to these changing domains. A highly effective CTTA method involves applying layer-wise adaptive learning rates for selectively adapting pre-trained layers. However, it suffers from the poor estimation of domain shift and the inaccuracies arising from the pseudo-labels. This work aims to overcome these limitations by identifying layers for adaptation via quantifying model prediction uncertainty without relying on pseudo-labels. We utilize the magnitude of gradients as a metric, calculated by backpropagating the KL divergence between the softmax output and a uniform distribution, to select layers for further adaptation. Subsequently, for the parameters exclusively belonging to these selected layers, with the remaining ones frozen, we evaluate their sensitivity to approximate the domain shift and adjust their learning rates accordingly. We conduct extensive image classification experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C, demonstrating the superior efficacy of our method compared to prior approaches.
Codehttps://github.com/sarthaxxxxx/PALM
@@ 1 Introduction
In a real-world setting, machine-perception systems (Arnold et al. 2019) function in a rapidly changing environment. Pre-trained vision models are susceptible to performance degradation caused by potential distribution shifts in such contexts. For instance, while these models may have been trained on clean images, images captured by sensors could be corrupted by various weather conditions.
In the realm of addressing domain shifts due to image corruptions, test-time adaptation (TTA) has been gaining a lot of traction lately (Döbler, Marsden, and Yang 2023; Wang et al. 2020). Here, a pre-trained source model, with no access to its training (source) data due to privacy, is adapted to unlabeled test data through online training. One strategy in TTA involves mitigating domain shifts by adjusting the source model parameters using pseudo labels and minimizing the entropy of the model predictions (Wang et al. 2020). While this approach effectively adapts the model to a single type of corruption, it faces two constraints when continually applied across a sequence of test tasks featuring different corruptions. Firstly, there is the issue of error accumulation since pseudo labels are generated for unlabeled test data, errors accumulate over time. Secondly, catastrophic forgetting occurs, resulting in a long-term loss of the model’s source knowledge due to adaptations and large parameter updates across tasks.
Continual test-time adaptation (CTTA) has been recently introduced to tackle these challenges. In (Niu et al. 2022; Gong et al. 2022; Wang et al. 2020; Schneider et al. 2020), stable optimization is achieved by updating the batch normalization (BN) parameters or re-formulating its statistics. CoTTA (Wang et al. 2022) includes full model updates, with a teacher-student network setup, to handle the above issues. To further improve the performance of CoTTA, EcoTTA (Song et al. 2023) introduces the attachment of “meta networks”, i.e., a BN layer and a convolution block, to the frozen source model for efficient adaptation.
More recently, the authors in (Lee et al. 2022) demonstrated that based on the type of distributional shift of the test data, a certain subset of layers of a pre-trained model can be “surgically” or heuristically chosen to fine-tune. Inspired by this finding, Layer-wise Auto-Weighting (LAW) (Park et al. 2024) was introduced recently. LAW utilizes the Fisher Information Matrix (FIM) (Sagun et al. 2017) to estimate domain shifts, automatically weighting trainable parameters and adjusting their learning rates (LRs) accordingly. While LAW achieves state-of-the-art results for CTTA, it still encounters two significant limitations: 1) It presupposes that layer-wise importance can be accurately estimated via FIM with initially generated pseudo-labels. Nevertheless, it is widely recognized that pseudo-labels are inherently noisy and unreliable. 2) Relying on directly accumulating such approximated importance from noisy labels to calculate domain shifts can lead to inaccuracies.
@@ 2 Related Works
Test-time adaptation (TTA): The goal of TTA is to adapt source models to unlabelled test data in real-time using a source-free online approach (Jain and Learned-Miller 2011; Sun et al. 2020; Wang et al. 2020). TENT (Wang et al. 2020) optimizes BN parameters by minimizing the Shannon entropy (Shannon 1948) of pseudo labels. BN Stats Adapt (BN-1) (Schneider et al. 2020) replaces the BN statistics with those from corrupted test data, while AdaContrast (Chen et al. 2022) uses contrastive learning and a memory module to update the entire model. However, these methods address only single domain shifts and overlook continual shifts.
Continual TTA (CTTA): CTTA requires models to adapt to dynamic domain shifts. CoTTA (Wang et al. 2022) uses a teacher-student setup with full model updates. The teacher model is updated using an exponential moving average (EMA) to deal with catastrophic forgetting and error accumulation with time. EcoTTA (Song et al. 2023) adds “meta networks” to a frozen source model, computing an $L_{1}$ regularization loss to prevent catastrophic forgetting. DSS (Wang et al. 2024) introduces a dynamic thresholding framework to differentiate between low and high sample quality.
Efficient model fine-tuning: Freezing model parameters to learn information from downstream data has witnessed a good body of work (Sener et al. 2016; Long et al. 2016; Zintgraf et al. 2019; Guo et al. 2019). Depending upon the distribution shift of the target task, in (Lee et al. 2022), the authors demonstrate that heuristically selecting a subset of convolutional blocks for training while freezing the remaining blocks works better than full fine-tuning.
Adaptive LR methods: Research such as (Yosinski et al. 2014; Zeiler and Fergus 2014) demonstrates that deeper layers of a network extract task-specific features, unlike shallow layers. AutoLR (Ro and Choi 2021) developed an adaptive LR scheme based on weight variations between layers, although these rates tend to mostly stay constant during training. For CTTA, LAW (Park et al. 2024) uses the FIM (Sagun et al. 2017) to assess domain shifts and adjust parameter LRs accordingly. However, using pseudo-labels to calculate FIM may lead to inaccuracies, and accumulating FIMs across domains can result in errors. Moreover, the adjustment of LR weights for each layer is constrained by the base LR, disregarding the relative “sensitivity” of parameters to the adaptation needs of the current batch and domain.
Parameter sensitivity: Neural network pruning literature extensively explores parameter sensitivity (LeCun, Denker, and Solla 1989; Luo, Wu, and Lin 2017; Theis et al. 2018; Molchanov et al. 2019). It suggests that less sensitive parameters require higher LRs for effective adaptation, and vice-versa. Additionally, (Liang et al. 2022) introduces a concept of temporal sensitivity variation, influenced by uncertainty, as a factor in training models for tasks like natural language understanding and machine translation.
@@ 3 Preliminary
Problem setup: CTTA aims to adapt a discriminative source model, parameterized as $\theta^{s}$ and trained on the source dataset $(\mathcal{X}^{\text{source}},\mathcal{Y}^{\text{source}})$, to multiple incoming tasks $\mathcal{T}_{i}=\{x_{k}\}_{k=1}^{K}$, at test-time. Each $\mathcal{T}_{i}$ signifies the $i^{th}$ task with $K$ batches, across varying data distributions. At time step t=0, the model is initialized to ${\theta^{s}}$. It is then gradually adapted to each incoming batch $x_{k}$ of $\mathcal{T}_{i}$ in an online manner, where the model parameters are updated to ${\theta^{t}}$. The source dataset is unavailable in this adaptation process due to privacy and storage constraints. With the continuum of tasks, at each $t$, the complete model parameters $\theta^{t}$ consist of a deep tensor with $N$ layers, expressed as $\theta^{t}=\{\theta^{t}_{1},\ldots,\theta^{t}_{N}\}$. Each layer $\theta^{t}_{n}$ can represented by $\theta^{t}_{n}=\{\theta^{t}_{1,n},\ldots,\theta^{t}_{j,n},\ldots\}$, where $\theta_{j,n}^{t}$ is an arbitrary parameter of layer $\theta_{n}^{t}$.
Core motivations: We provide a detailed view of surgical fine-tuning (Lee et al. 2022), AutoLR (Ro and Choi 2021), and LAW (Park et al. 2024). It is critical to understand the design principles and reasons for having an adaptive LR scheme for CTTA, which motivates this work.
@@
Inspired by (Yosinski et al. 2014; Zeiler and Fergus 2014), AutoLR, proposed for supervised model fine-tuning on a downstream task, claims that the weight changes across layers should be in an increasing order, where the weight change $\delta w_{n}$ in the $n^{th}$ layer with weights $w_{n}$ at training epoch e is $\delta w_{n}=\|w_{n}^{e}\ -w_{n}^{(e-1)}\|$. They hypothesize that initial layers, which extract general features, should have minimal $\delta w_{n}$, while deeper layers extract task-specific features. An adaptive LR optimization scheme constrains weight variations to increase across layers, without assessing each layer’s importance during adaptation. However, the LRs for each layer show minimal deviation and remain consistent throughout the training process.
As the first adaptive LR work for CTTA, LAW hypothesizes that identifying and controlling the adaptation rate of the different layers is useful (Lee et al. 2022). To understand the loss landscape and layer importance, the authors compute the gradient from the log-likelihood as $s(\theta_{j,n}^{t};x_{k})=\nabla_{\theta_{j,n}^{t}}\log(p(x_{k}))$, where $p(x_{k})$ is the predicted output. Then, the FIM (Sagun et al. 2017), as a domain-shift indicator, is computed as a first-order approximation of the Hessian matrix. This layer-wise importance scheme completely relies on the pseudo-labels to compute the log-likelihood and is highly unreliable and uncertain. Due to the distribution shift, the label quality would begin to decrease for a continuum of tasks. The authors also propose to capture “domain-level” FIM ($\hat{F_{n}^{t}}$), for each layer, by a direct accumulation of the current layer-wise FIM (${F_{n}^{t}}$) as, $\hat{F}_{n}^{t}=\hat{F}_{n}^{t-1}+F_{n}^{t}$.
This means that LAW puts equal emphasis on $\hat{F}_{n}^{t}$ and ${F_{n}^{t}}$, for each layer. An issue arises when ${F}_{n}^{t}$ of the current batch is small but gets outweighed by a “relatively” larger $\hat{F}_{n}^{t}$. This raises questions on the amount of adaptation required by the $n^{th}$ layer for batch $x_{k}$ as the FIMs dynamically change for each test batch. This leads to a possible imbalance between the accumulated FIMs $\hat{F}_{n}^{t}$ and the current FIM ${F_{n}^{t}}$.
@@
4 Proposed Methodology
Figure 1: The framework of our proposed method PALM.
Figure 1 illustrates our approach in two stages.
Layer selection by quantifying prediction uncertainty: CTTA maximizes model performance in the current domain and so it is vital to assess prediction certainty amidst rapid domain shifts. In the literature on continual learning and TTA, several works utilize FIM (Kirkpatrick et al. 2017; Schwarz et al. 2018; Park et al. 2024). However, the unavailability of true labels, with unknown distribution, pushes us to capture the model prediction uncertainty and adapt selected layers, automatically chosen, to a batch of test data, unlike surgical fine-tuning (Lee et al. 2022).
Motivated by (Huang, Geng, and Li 2021), which proposed an uncertainty scoring function from the gradient space, we compute the gradients of each layer. In essence, we backpropagate the KL divergence between the softmax probabilities and a uniform distribution u = [$\frac{1}{C}$, …, $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where $C$ is the total number of classes of the current task. This is a measurement of how far the distribution of the model’s prediction is from a uniform distribution. Intuitively, the KL divergence will be small with a large distribution shift. Let $h(\cdot)$ denote the classifier of the model. The loss $\mathcal{L}(\theta^{t})$, to compute the gradients, is computed as follows,
@@ 4 Proposed Methodology
Parameter sensitivity as a domain shift indicator: New studies have found that model parameters demonstrate levels of criticality (Chatterji, Neyshabur, and Sedghi 2019; Zhang, Bengio, and Singer 2022) across different positions. Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors (Mozer and Smolensky 1988; Rasmussen and Ghahramani 2000). Inspired by this observation, we assert that parameters contribute differently during CTTA. From the previous stage, we automatically select the layers based on the prediction uncertainty. For parameters of these selected layers only, we compute their sensitivity i.e., an approximation of an error when it is removed. Motivated by (LeCun, Denker, and Solla 1989; Molchanov et al. 2016, 2019), we compute the sensitivity of a parameter $\theta_{j,n}^{t}$ based on the loss (Eq. 1) as follows,
@@
Precisely, $S_{j,n}^{t}$ approximates the change in the loss landscape at $\theta^{t}$ when $\theta_{j,n}^{t}$ is removed. Sensitivity, computed as a “weighting factor” of model weights (Molchanov et al. 2016, 2019; Liang et al. 2022), determines the level of adaptation needed for model parameters. We utilize the gradients of parameters from the selected layers, computed by Eq. 2, to compactly derive Eq. 3 from the first-order Taylor series expansion of $\mathcal{L}(\theta^{t})$ (Lee, Ajanthan, and Torr 2018) to,
@@
Our rationale is as follows: parameters with low sensitivity are pushed to have higher LRs owing to more adaptation. However, $S_{j,n}^{t}$ of $\theta_{j,n}^{t}$ can be highly uncertain and would have a large variation. We introduce a method to capture this “domain-level” uncertainty by computing a weighted moving average of the sensitivity $S_{j,n}^{t}$. This is formulated as,
@@
where $\hat{S}_{j,n}^{t}$ denotes the “domain-level” sensitivity, computed on-the-fly, and $\alpha$in [0,1] is a smoothing factor. Specifically, $\alpha$ decides the contribution of the current sensitivity $S_{j,n}^{t}$ to $\hat{S}_{j,n}^{t}$. Since CTTA happens at the batch level, it is critical to achieve a good balance between $S_{j,n}^{t}$ and $\hat{S}_{j,n}^{t}$, with a slightly larger focus on $S_{j,n}^{t}$. As can be seen, we do not directly accumulate the “domain-level” information, unlike LAW. Due to large distribution shifts happening because of the continuum of tasks, larger deviations of $S_{n}^{t}$ from $\hat{S}_{n}^{t}$ are bound to happen. We quantify this statistic as an absolute difference,
@@
We emphasize that since Eq. 5 leverages sensitivity from previously computed batches and domains, $\hat{D}_{j,n}^{t}$ would act as a good indicator of how uncertain $S_{j,n}^{t}$ is to the previous batches and domains i.e. $\hat{S}_{j,n}^{t}$, for $\theta_{j,n}^{t}$. This would take care of error accumulation too. To automatically alter the LRs of these parameters, the “importance” ($i_{j,n}^{t}$) is calculated as,
@@
Optimization: Though our method is robust, it is necessary to avoid miscalibrated labels. TENT (Wang et al. 2020) minimized the entropy on the output logits, l, following $\mathcal{H}(l)=-\sum_{c}p(l_{c})\log(p(l_{c}))$, where $p(l_{c})$ is the probability for class c. Following EcoTTA (Song et al. 2023), we use entropy minimization as the objective for adaptation where $\mathcal{H}(l)$ is minimized for samples below a certain fixed threshold $\mathcal{H}_{0}$ as,
@@
$\mathcal{H}_{0}$ is set to $0.4\times\log C$ following EcoTTA. For a fair comparison with prior works (Wang et al. 2022; Park et al. 2024), we apply a consistency loss between l and $\hat{l}$, computed as the output logits from the augmented version of $x_{k}$, to regularize the model as follows,
@@
5 Experiments and Results